OpenStudy (anonymous):

Rewrite y = x2 + 14x + 29 in general form.

OpenStudy (anonymous):

*y = x^2 +14 + 29

OpenStudy (amistre64):

which form is general?

OpenStudy (anonymous):

What do you mean by general form? Amistre surely knows

OpenStudy (anonymous):

maybe not :-)

OpenStudy (amistre64):

terminology is not my strong suite :)

OpenStudy (anonymous):

I'm not sure. It just says in general. The name of the lesson is "Completing the Square".. I didn't learn it yet though.

OpenStudy (anonymous):

Ok completing the square: y=x^2 + 14x + 29=(x+7)^2-49+29=(x+7)^2-20